0=xx(squared)-50x+175 ????? I tried to use the quadratic formula, but it became a imaginary number, and I need to have a positive number.
thanks
-tiproggrammer

Answer by stanbon(75887)

0=xx(squared)-50x+175
----------
x^2-50x+175=0
x = [50 +- sqrt(2500-4*175)]/2
x = [50 +- sqrt1800)]/2
x = [50 +- 30sqrt2]/2
x = 25 + 15sqrt2 or x = 25 - 15sqrt2
===============
Cheers,
Stan H.